A 50 kg object accelerates uniformly from rest to a speed of 10.0 m/s in 0.5 seconds. The object's acceleration is ___.
1
Expert's answer
2020-12-09T09:43:35-0500

By definition, the acceleration is:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 10m\/s" is the final velocity, "v_i = 0m\/s" is the initial velocity (started from rest), and "t=0.5s" is the the time interval. Thus, the acceleration will be:


"a = \dfrac{10m/s-0m/s}{0.5} = 20m/s^2"

Answer. 20 m/s^2.